given a rectangle whose width, w, is 5 less than twice the length, L, state the perimeter of the rectangle as a function of L.
Answer by mananth(16946)

w width
w= 2L-5
...
Perimeter = 2*(L+W)
P= 2(L+(2l-5)
P=2(3L-5)
P=6L-10